y = log π is equivalent to:
a. 10^-y = π
b. 10^π = y
c. 10^y = π
d. Cannot be defined without first knowing the base of the logarithm

Found 2 solutions by stanbon, Nate:
Answer by stanbon(75887)

Ans: c
Cheers,
Stan H.

Answer by Nate(3500)

y = log π
10^y = 10^log π
* 10^log π = π
10^y = π
c.)